PARALLEL CIRCUIT


  1. what is the equivalent resistance of this circuit?

(150 ohms upper side of the circuit, 250 ohms in the left side and 50 ohms below side of the circuit)


1
Expert's answer
2021-05-24T14:03:52-0400

Since, "150 \Omega", "250 \Omega", and "50 \Omega" are in parallel.

Equivalent of the resistances in parallel is given by,

"\frac{1}{R} = \frac{1}{R_1} + \frac{1}{R_2}+\frac{1}{R_3}"


Putting values,

"\frac{1}{R} = \frac{1}{150} + \frac{1}{250}+\frac{1}{50} = \frac{5+3+15}{750}= \frac{23}{750}"



"R = \frac{750}{23} = 32.61 \Omega"